# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Acknowledgements.
S2.1.2 Simulated Annealing
Description
@@
Numerical Implementation
We test both linear and geometric temperature schedules, controlled by the number of steps $k$. The initial hot temperature is noted $T_{\textrm{hot}}\equiv T_{\ell=1}$, the final cold temperature is noted $T_{\textrm{cold}}\equiv T_{\ell=k}$, and intermediate temperatures are noted by $T_{\ell}$. In the linear interpolation scheme, the intermediate temperatures are defined as
$T^{-1}_{\ell}=T_{\textrm{hot}}^{-1}+\ell\frac{T_{\textrm{cold}}^{-1}-T_{\textrm{hot}}^{-1}}{k}\quad\textrm{where}\quad\ell=1,~{}2,~{}\ldots,~{}k.$ (S5)
In the geometric interpolation scheme, the intermediate temperatures are defined as
$T^{-1}_{\ell}=\exp\Biggl{(}\ln T_{\textrm{hot}}^{-1}+\ell\frac{\ln T_{\textrm{cold}}^{-1}-\ln T_{\textrm{hot}}^{-1}}{k}\Biggr{)}\quad\textrm{where}\quad\ell=1,~{}2,~{}\ldots,~{}k.$ (S6)
Each one of these $k$ steps involves one sweep, which is an update on every single spin of the model. We choose
$T_{\textrm{hot}}=6/\ln 2\quad\textrm{and}\quad T_{\textrm{cold}}=2/\ln\bigl{(}100N\bigr{)}.$ (S7)
The rationale behind those numbers is as follows. At first, we want fast mixing where all spins have a high probability of flipping. The three-fold connectivity of the $3$-regular graphs with $\textrm{max}(\mathsf{W}_{ij})=1$ is such that the largest energy difference between two configurations is $\tilde{C}(\boldsymbol{z})-\tilde{C}(\boldsymbol{z^{\prime}})=6$. We want such a transition to be accepted half of the time. Solving Eq. (S4) with those numbers leads to $T_{\textrm{hot}}=6/\ln 2$. Toward the end, in the last sweeps, the state should have a low probability of being excited as it should correspond to a local/global minimum. For unit-weight $3$-regular graphs, the smallest excitation, or smallest energy difference between two configurations is $\tilde{C}(\boldsymbol{z})-\tilde{C}(\boldsymbol{z^{\prime}})=2$. We bound the probability to excite any of the $N$ spins by $1\%$, leading to the equation $1\%=Ne^{2/T_{\textrm{cold}}}$, which we solve for $T_{\textrm{cold}}=2/\ln(100N)$.
We report and discuss the relative performance of the linear and geometric temperature interpolation schemes in Sec. S5.2. We find that the geometric temperature schedule yields better results for the cases and regimes considered in this work.
Algorithmic Complexity
@@
S2.1.3 Parallel Tempering
Description
@@
$p_{\textrm{accept}}\bigl{(}\boldsymbol{z}_{\ell}\leftarrow\boldsymbol{z}_{\ell+1}\bigr{)}=\textrm{min}\Biggl{\{}1,e^{\bigl{[}\tilde{C}(\boldsymbol{z}_{\ell})-\tilde{C}(\boldsymbol{z}_{\ell+1})\bigr{]}\bigl{[}T^{-1}_{\ell}-T^{-1}_{\ell+1}\bigr{]}}\Biggr{\}}.$ (S8)
@@
Numerical Implementation
We employ the same cold and hot temperatures that were defined for simulated annealing in Eq. (S7), based on the properties of unit-weight $3$-regular graphs. We fix the number of replicas to $n=10$ defined from $T_{\textrm{hot}}$ to $T_{\textrm{cold}}$ with intermediate temperatures defined using a geometric spacing. We perform one sweep on each individual replica before performing the replica exchange step, starting from the highest temperature $T_{\textrm{hot}}$.
Algorithmic Complexity
@@
S2.1.4 The Burer-Monteiro Solver: Nonlinear Optimization Heuristic with Local Search
Description
The Burer-Monteiro solver is based on a relax-and-round scheme of the maximum cut objective function [15, 16]. The relaxation is such that the resulting optimization problem is nonlinear with a nonconvex objective function. The nonconvexity may sound daunting since it means that, in general, the objective function has multiple local nonglobal minima. In practice, the Burer-Monteiro algorithm is found to return high-quality solutions, making it one of the best heuristics to date for the maximum cut problem [56].
Numerical Implementation
We rely on a numerical implementation of the Burer-Monteiro solver which is part of the MQLib library [56] under the name ‘Burer2002’. The fundamental control parameter of the Burer-Monteiro algorithm is its number of steps, corresponding to the number of random perturbations explored from the current minimum. However, the MQLib implementation only exposes a runtime limit where a number of steps will be performed until the runtime exceeds the limit. The algorithm always runs for at least one full step.
Algorithmic Complexity
@@
S2.1.5 The Commercial Solver Gurobi
Description
Since Gurobi [13] cannot directly handle an Ising model in the form of Eq. (S1), where variable $z_{i}$ takes values in {±$\{\pm 1\}$, we first reformulate it into a quadratic unconstrained binary optimization problem (QUBO) to minimize
$\tilde{C}_{\textrm{QUBO}}(\boldsymbol{x})=\frac{1}{2}\sum\nolimits_{i,j=1}^{N}\mathsf{W}_{ij}\bigl{(}2x_{i}-1\bigr{)}\bigl{(}2x_{j}-1\bigr{)},$ (S9)
@@
We employ Gurobi 11.0.0, with the runtime limit set to $600$ seconds and the NonConvex parameter set to $2$, to solve nonconvex quadratic problems by translating them into bilinear form and applying spatial branching. All tests ran on a Linux cluster with 48 AMD EPYC 7643 2.3GHz CPUs and 1 TB RAM, with each test restricted to using only a single thread.
Algorithmic Complexity
@@
Analyzing the complexity of the branch-and-cut method in Gurobi is challenging and is affected by branching, node selection, cut generation, and heuristic strategies. Generally, the time complexity of the branch-and-bound algorithm is measured as $O(b^{d})$, where $b$ represents the branching factor, and $d$, the solution’s depth, is equivalent to $N$ in the $3$-regular graphs problems of Eq. (S1). The space complexity is typically $O(bd)$, reflecting the maximum number of nodes stored in memory at any point. Despite the potential worst-case time complexity, the branch-and-cut method proves efficient in practice by using effective bounding techniques to prune a large section of the tree.
S2.1.6 Greedy Solver
Description
@@
We employ a classical randomized greedy solver [24]. It builds a solution $\boldsymbol{z}\in\{\pm 1\}^{N}$ iteratively. The vector is initially set to zero, i.e., $\boldsymbol{z}=\boldsymbol{0}$, and the algorithm runs as follows. One selects uniformly, at random, a variable $z_{i}$. One sets this variable to $+1$ and computes the corresponding objective value $\tilde{C}(\boldsymbol{z})$ according to Eq. (S1). One repeats this step by now setting the variable to $-1$. Finally, the variable is set once and for all to the value $z_{i}=\pm 1$ extremizing the objective function. Ties are broken down uniformly at random. The steps are repeated until all $N$ variables are set to ±$\pm 1$.
Algorithmic Complexity
@@
The $N$-variable solution is built over $N$ iterations. Each of these iterations requires evaluating the objective function $\tilde{C}$ of Eq. (S1) twice. At each iteration, one only needs to compute terms in $C$ involving the selected variable. For $3$-regular graph problems, there are only three such terms. Thus, the algorithmic complexity of evaluating the objective function is independent of $N$. As such, the global algorithmic complexity of the classical greedy solver is $O(N)$ for $3$-regular graph problems.
S2.1.7 Relax-and-Round Approach
Description
@@
This approach first relaxes the constraint on the solution space from $\boldsymbol{z}\in\{\pm 1\}^{N}$ to $\boldsymbol{z}\in\mathbb{R}^{N}$ with $\|\boldsymbol{z}\|=\textrm{constant}$. Thus, the extremization task of Eq. (S1) becomes an eigenvalue problem for which the leading eigenvector of $\mathsf{W}$ is the optimal solution. The leading eigenvector is then sign-rounded entrywise to recover a valid solution. In practice, we compute the $k=8$ leading eigenvectors. Each of them is sign-rounded, and the one yielding the best objective value is returned as the solution from the solver.
Algorithmic Complexity
@@
S2.1.8 Goemans-Williamson Algorithm
Description
@@
We employ a formulation of the Goemans-Williamson algorithm [12, 60] based on an eigenvalue problem [61, 62, 63, 64] due to the quantum analog that we developed. The Goemans-Williamson algorithm seeks to solve the maximum cut problem, corresponding to maximizing the objective function
$C(\boldsymbol{z})=\frac{1}{4}\sum\nolimits_{i,j=1}^{N}\mathsf{W}_{ij}\Bigl{(}1-z_{i}z_{j}\Bigr{)},$ (S10)
@@
as defined in the main text. From a practical point of view, the tasks of maximizing Eq. (S10) and minimizing Eq. (S1) lead to the same optimal solution $\boldsymbol{z}_{\textrm{opt}}\in\{\pm 1\}^{N}$, albeit with a different objective value $\tilde{C}(\boldsymbol{z}_{\textrm{opt}})\neq C(\boldsymbol{z}_{\textrm{opt}})$. Therefore, one can use the Goemans-Williamson algorithm for solving the original problem of interest in Eq. (S1), which has been rewritten in the form of Eq. (S10).
@@
We introduce the Laplacian matrix $\mathsf{L}=\mathsf{D}-\mathsf{W}$ where $\mathsf{D}_{ij}=\delta_{ij}\sum_{k}\mathsf{W}_{ik}$ is the degree matrix with $\delta_{ij}$ the Kronecker delta. For unit-weight $3$-regular graph problems, $\mathsf{D}=3\mathsf{I}$, where $\mathsf{I}$ is the identity matrix. Thus, the objective function of Eq. (S10) can be written as $C(\boldsymbol{z})=\sum\nolimits_{i,j=1}^{N}\mathsf{L}_{ij}z_{i}z_{j}/4$. First, the largest algebraic eigenvalue of $N\mathsf{L}/4$ is an upper bound to the optimal objective value $C(\boldsymbol{z}_{\textrm{opt}})$ [61, 62, 63, 64]. Second, one notes that the objective function is invariant under the transformation $\mathsf{L}\to\mathsf{L}+\textrm{diag}(u)$, where $\textrm{diag}(u)$ is a traceless diagonal matrix formed by the vector $\boldsymbol{u}\in\mathbb{R}^{N}$, known as the correcting vector. This means that the optimal objective value $C(\boldsymbol{z}_{\textrm{opt}})$ is also bounded by the leading eigenvalue of $N[\mathsf{L}+\textrm{diag}(u)]/4$, given the constraint $\sum_{i}u_{i}=0$. Hence, one can leverage the additional degrees of freedom introduced by the correcting vector $\boldsymbol{u}$ to tighten the eigenvalue bound. In practice, we solve a relaxed version, i.e., $\boldsymbol{z}\in\{\pm 1\}^{N}\to\boldsymbol{z}\in\mathbb{R}^{N}$
$\textrm{min}_{\sum_{i}u_{i}=0}~{}\textrm{max}_{\|\boldsymbol{z}\|=1}\boldsymbol{z}^{T}\frac{N}{4}\Bigl{[}\mathsf{D}-\mathsf{W}+\textrm{diag}(u)\Bigr{]}\boldsymbol{z},$ (S11)
@@
and then, sign-round the leading eigenvector, i.e., $\boldsymbol{z}\in\mathbb{R}^{N}\to\boldsymbol{z}\in\{\pm 1\}^{N}$, to recover a valid solution with respect to the original problem. This algorithm is equivalent to the celebrated Goemans-Williamson algorithm for the maximum cut problem [61, 62, 63, 64, 12, 60].
Remarks
@@
For unit-weight $3$-regular graph problems, the degree matrix is a constant diagonal matrix. Therefore, it is irrelevant for finding the leading eigenvector of Eq. (S11). Now, if one sets $\boldsymbol{u}=\boldsymbol{0}$, the method maps to the relax-and-round approach developed in Sec. S2.1.7. As such, it may seem like the relax-and-round approach of Sec. S2.1.7 gives a lower-bound to the solution of the Goemans-Williamson algorithm. However, the relevant solution is the sign-rounded one, which might not be better in the $\boldsymbol{u}$-augmented version.

### v2

@@ B.1.1 Markov Chain Monte-Carlo
B.1.2 Simulated Annealing
Description—
@@ B.1.2 Simulated Annealing
Numerical Implementation—
We test both linear and geometric temperature schedules, controlled by the number of sweeps $k$ (at each sweep, the temperature is reduced). The initial hot temperature is noted $T_{\textrm{hot}}\equiv T_{\ell=1}$, the final cold temperature is noted $T_{\textrm{cold}}\equiv T_{\ell=k}$, and intermediate temperatures are noted by $T_{\ell}$. In the linear interpolation scheme, the intermediate temperatures are defined as
$T^{-1}_{\ell}=T_{\textrm{hot}}^{-1}+\ell\frac{T_{\textrm{cold}}^{-1}-T_{\textrm{hot}}^{-1}}{k}$ (8)
where $\ell=1,~{}2,~{}\ldots,~{}k$. In the geometric interpolation scheme, the intermediate temperatures are defined as
$T^{-1}_{\ell}=\exp\Biggl{(}\ln T_{\textrm{hot}}^{-1}+\ell\frac{\ln T_{\textrm{cold}}^{-1}-\ln T_{\textrm{hot}}^{-1}}{k}\Biggr{)}$ (9)
where $\ell=1,~{}2,~{}\ldots,~{}k$. We choose
$T_{\textrm{hot}}=6/\ln 2\quad\textrm{and}\quad T_{\textrm{cold}}=2/\ln\bigl{(}100N\bigr{)}.$ (10)
The rationale behind those numbers is as follows. At first, we want fast mixing where all spins have a high probability of flipping. The three-fold connectivity of the $3$-regular graphs with $\textrm{max}(\mathsf{W}_{ij})=1$ is such that the largest energy difference between two configurations is $\tilde{C}(\boldsymbol{z})-\tilde{C}(\boldsymbol{z^{\prime}})=6$. We want such a transition to be accepted half of the time. Solving Eq. (7) with those numbers leads to $T_{\textrm{hot}}=6/\ln 2$. Toward the end, in the last sweeps, the state should have a low probability of being excited as it should correspond to a local/global minimum. For unit-weight $3$-regular graphs, the smallest excitation, or smallest energy difference between two configurations is $\tilde{C}(\boldsymbol{z})-\tilde{C}(\boldsymbol{z^{\prime}})=2$. We bound the probability to excite any of the $N$ spins by $1\%$, leading to the equation $1\%=Ne^{2/T_{\textrm{cold}}}$, which we solve for $T_{\textrm{cold}}=2/\ln(100N)$.
We report and discuss the relative performance of the linear and geometric temperature interpolation schemes in App. E.2. We find that the geometric temperature schedule yields better results for the cases and regimes considered in this work.
Algorithmic Complexity—
@@
B.1.3 Parallel Tempering
Description—
@@ B.1.3 Parallel Tempering
$p_{\textrm{accept}}\bigl{(}\boldsymbol{z}_{\ell}\leftarrow\boldsymbol{z}_{\ell+1}\bigr{)}=\textrm{min}\Biggl{\{}1,e^{\bigl{[}\tilde{C}(\boldsymbol{z}_{\ell})-\tilde{C}(\boldsymbol{z}_{\ell+1})\bigr{]}\bigl{[}T^{-1}_{\ell}-T^{-1}_{\ell+1}\bigr{]}}\Biggr{\}}.$ (11)
@@
Numerical Implementation—
We employ the same cold and hot temperatures that were defined for simulated annealing in Eq. (10), based on the properties of unit-weight $3$-regular graphs. We fix the number of replicas to $n=10$ defined from $T_{\textrm{hot}}$ to $T_{\textrm{cold}}$ with intermediate temperatures defined using a geometric spacing. We perform one sweep on each individual replica before performing the replica exchange step, starting from the highest temperature $T_{\textrm{hot}}$.
Algorithmic Complexity—
@@
B.1.4 The Burer-Monteiro Solver: Nonlinear Optimization Heuristic with Local Search
Description—
The Burer-Monteiro solver is based on a relax-and-round scheme of the maximum cut objective function [16, 17]. The relaxation is such that the resulting optimization problem is nonlinear with a nonconvex objective function. The nonconvexity may sound daunting since it means that, in general, the objective function has multiple local nonglobal minima. In practice, the Burer-Monteiro algorithm is found to return high-quality solutions, making it one of the best heuristics to date for the maximum-cut problem [54].
Numerical Implementation—
We rely on a numerical implementation of the Burer-Monteiro solver which is part of the MQLib library [54] under the name ‘Burer2002’. The fundamental control parameter of the Burer-Monteiro algorithm is its number of steps, corresponding to the number of random perturbations explored from the current minimum. However, the MQLib implementation only exposes a run-time limit where a number of steps will be performed until the run-time exceeds the limit. The algorithm always runs for at least one full step.
Algorithmic Complexity—
@@ B.1.4 The Burer-Monteiro Solver: Nonlinear Optimization Heuristic with Local Search
B.1.5 The Commercial Solver Gurobi
Description—
Since Gurobi [14] cannot directly handle an Ising model in the form of Eq. (4), where variable $z_{i}$ takes values in {±$\{\pm 1\}$, we first reformulate it into a quadratic unconstrained binary optimization problem (QUBO) to minimize
$\tilde{C}_{\textrm{QUBO}}(\boldsymbol{x})=\frac{1}{2}\sum\nolimits_{i,j=1}^{N}\mathsf{W}_{ij}\bigl{(}2x_{i}-1\bigr{)}\bigl{(}2x_{j}-1\bigr{)},$ (12)
@@ B.1.5 The Commercial Solver Gurobi
We employ Gurobi 11.0.0, with the run-time limit set to $600$ seconds and the NonConvex parameter set to $2$, to solve nonconvex quadratic problems by translating them into bilinear form and applying spatial branching. All tests ran on a Linux cluster with 48 AMD EPYC 7643 2.3GHz CPUs and 1 TB RAM, with each test restricted to using only a single thread.
Algorithmic Complexity—
@@
Analyzing the complexity of the branch-and-cut method in Gurobi is challenging and is affected by branching, node selection, cut generation, and heuristic strategies. Generally, the time complexity of the branch-and-bound algorithm is measured as $O(b^{d})$, where $b$ represents the branching factor, and $d$, the solution’s depth, is equivalent to $N$ in the $3$-regular graphs problems of Eq. (4). The space complexity is typically $O(bd)$, reflecting the maximum number of nodes stored in memory at any point. Despite the potential worst-case time complexity, the branch-and-cut method proves efficient in practice by using effective bounding techniques to prune a large section of the tree.
B.1.6 Greedy Solver
Description—
@@ B.1.6 Greedy Solver
We employ a classical randomized greedy solver [58]. It builds a solution $\boldsymbol{z}\in\{\pm 1\}^{N}$ iteratively. The vector is initially set to zero, i.e., $\boldsymbol{z}=\boldsymbol{0}$, and the algorithm runs as follows. One selects uniformly, at random, a variable $z_{i}$. One sets this variable to $+1$ and computes the corresponding objective value $\tilde{C}(\boldsymbol{z})$ according to Eq. (4). One repeats this step by now setting the variable to $-1$. Finally, the variable is set once and for all to the value $z_{i}=\pm 1$ extremizing the objective function. Ties are broken down uniformly at random. The steps are repeated until all $N$ variables are set to ±$\pm 1$.
Algorithmic Complexity—
@@
The $N$-variable solution is built over $N$ iterations. Each of these iterations requires evaluating the objective function $\tilde{C}$ of Eq. (4) twice. At each iteration, one only needs to compute terms in $C$ involving the selected variable. For $3$-regular graph problems, there are only three such terms. Thus, the algorithmic complexity of evaluating the objective function is independent of $N$. As such, the global algorithmic complexity of the classical greedy solver is $O(N)$ for $3$-regular graph problems.
B.1.7 Relax-and-Round Approach
Description—
@@ B.1.7 Relax-and-Round Approach
This approach first relaxes the constraint on the solution space from $\boldsymbol{z}\in\{\pm 1\}^{N}$ to $\boldsymbol{z}\in\mathbb{R}^{N}$ with $\|\boldsymbol{z}\|=\textrm{constant}$. Thus, the extremization task of Eq. (4) becomes an eigenvalue problem for which the leading eigenvector of $\mathsf{W}$ is the optimal solution. The leading eigenvector is then sign-rounded entrywise to recover a valid solution. In practice, we compute the $k=8$ leading eigenvectors. Each of them is sign-rounded, and the one yielding the best objective value is returned as the solution from the solver.
Algorithmic Complexity—
@@
B.1.8 Goemans-Williamson Algorithm
Description—
@@ B.1.8 Goemans-Williamson Algorithm
We employ a formulation of the Goemans-Williamson algorithm [13, 59] based on an eigenvalue problem [60, 61, 62, 63] due to the quantum analog that we developed. The Goemans-Williamson algorithm seeks to solve the maximum-cut problem, corresponding to maximizing the objective function
@@
as defined in the main text. From a practical point of view, the tasks of maximizing Eq. (13) and minimizing Eq. (4) lead to the same optimal solution $\boldsymbol{z}_{\textrm{opt}}\in\{\pm 1\}^{N}$, albeit with a different objective value $\tilde{C}(\boldsymbol{z}_{\textrm{opt}})\neq C(\boldsymbol{z}_{\textrm{opt}})$. Therefore, one can use the Goemans-Williamson algorithm for solving the original problem of interest in Eq. (4), which has been rewritten in the form of Eq. (13).
@@
We introduce the Laplacian matrix $\mathsf{L}=\mathsf{D}-\mathsf{W}$ where $\mathsf{D}_{ij}=\delta_{ij}\sum_{k}\mathsf{W}_{ik}$ is the degree matrix with $\delta_{ij}$ the Kronecker delta. For unit-weight $3$-regular graph problems, $\mathsf{D}=3\mathsf{I}$, where $\mathsf{I}$ is the identity matrix. Thus, the objective function of Eq. (13) can be written as $C(\boldsymbol{z})=\sum\nolimits_{i,j=1}^{N}\mathsf{L}_{ij}z_{i}z_{j}/4$. First, the largest algebraic eigenvalue of $N\mathsf{L}/4$ is an upper bound to the optimal objective value $C(\boldsymbol{z}_{\textrm{opt}})$ [60, 61, 62, 63]. Second, one notes that the objective function is invariant under the transformation $\mathsf{L}\to\mathsf{L}+\textrm{diag}(u)$, where $\textrm{diag}(u)$ is a traceless diagonal matrix formed by the vector $\boldsymbol{u}\in\mathbb{R}^{N}$, known as the correcting vector. This means that the optimal objective value $C(\boldsymbol{z}_{\textrm{opt}})$ is also bounded by the leading eigenvalue of $N[\mathsf{L}+\textrm{diag}(u)]/4$, given the constraint $\sum_{i}u_{i}=0$. Hence, one can leverage the additional degrees of freedom introduced by the correcting vector $\boldsymbol{u}$ to tighten the eigenvalue bound. In practice, we solve a relaxed version, i.e., $\boldsymbol{z}\in\{\pm 1\}^{N}\to\boldsymbol{z}\in\mathbb{R}^{N}$
$\textrm{min}_{\sum_{i}u_{i}=0}~{}\textrm{max}_{\|\boldsymbol{z}\|=1}\boldsymbol{z}^{T}\frac{N}{4}\Bigl{[}\mathsf{D}-\mathsf{W}+\textrm{diag}(u)\Bigr{]}\boldsymbol{z},$ (14)
@@
and then, sign-round the leading eigenvector, i.e., $\boldsymbol{z}\in\mathbb{R}^{N}\to\boldsymbol{z}\in\{\pm 1\}^{N}$, to recover a valid solution with respect to the original problem. This algorithm is equivalent to the celebrated Goemans-Williamson algorithm for the maximum-cut problem [60, 61, 62, 63, 13, 59].
Remarks—
@@
For unit-weight $3$-regular graph problems, the degree matrix is a constant diagonal matrix. Therefore, it is irrelevant for finding the leading eigenvector of Eq. (14). Now, if one sets $\boldsymbol{u}=\boldsymbol{0}$, the method maps to the relax-and-round approach developed in App. B.1.7. As such, it may seem like the relax-and-round approach of App. B.1.7 gives a lower-bound to the solution of the Goemans-Williamson algorithm. However, the relevant solution is the sign-rounded one, which might not be better in the $\boldsymbol{u}$-augmented version.